OpenStudy (anonymous):

MEDAL AND FAN ! what is the approximate length of a semicircle with a diameter of 64 meters? a.100.48m b.200.96m c.401.92m d.1607.68m

OpenStudy (anonymous):

the formula for the circumference or length of a semicircle is

OpenStudy (anonymous):

(pi)r+d

OpenStudy (anonymous):

so you just plug in the values and you get A. i am pretty sure its a

OpenStudy (jdoe0001):

http://mathb.in/12760

OpenStudy (triciaal):

D = 64 R = 1/2 D = 32 C = 2 PI R SEMI = 1/2 PI = 3.142 LENGTH OF 1/2 C =1/2(2)(3.142)(32)

OpenStudy (anonymous):

thx guys

OpenStudy (anonymous):

@triciaal but you have to add another diameter because you have to close the semicircle